Two charges located on the positive x-axis has a first charge of 5nc which is 2cm from the orgin and second charge of 9nc which is 4cm from the orgin, what is the magnitude of force exerted by these two force on a third charge 5nc located at the origin

1
Expert's answer
2022-03-14T13:06:27-0400

Force

"F_{net}=\frac{kq_1q_2}{r^2}-\frac{kq_3q_2}{r^2}"


"F_{net}=9\times10^{9}\times5\times10^{-9}(\frac{5\times10^{-9}}{0.04^2}-\frac{5\times10^{-9}}{0.04^2})"

"F_{net}=0N"